Louisiana State University Health Sciences Center Administration & Finance
 

From: Ronnie Smith, Vice Chancellor, Administration and Finance
Subject: Hurricane Katrina Recovery – Update #49
Posted: August 17, 2007

The contractor who is remodeling the MEB cafeteria is completing all construction items this weekend except for the custom built-in cabinets and installation of new custom furniture. LSUHSC Facility Services intends to relocate tables and chairs being used in Seminar Rooms 4 and 5 to the cafeteria on Sunday. Starting after the conclusion of lunch service on Friday, August 17th Seminar Rooms 4 and 5 will be cleaned and returned to their normal classroom configuration.

The renovated cafeteria will be open for breakfast/lunch service commencing at 7:00 AM on Monday, August 20, 2007. Hours of operation will remain 7:00 am to 3:00 pm for the next several weeks until all the final touches are complete. Expect to see continued improvements such as custom furniture, cabinetry and all new tables and chairs. Once all work is complete and the new furniture is installed, it will be open from 7:00 am to 3:00 pm Monday through Friday for food service and remain open 24/7 as a student lounge/study area. Look for the grand opening in September.

FEMA has completed their initial review of three Section 406 Hazard Mitigation design concepts that were submitted for the Allied Health/Nursing Building, the Residence Hall, and the Resource Center. The architects have been asked to move forward with the more detailed Design/Development proposals for each building and they plan to deliver these in September. Once the internal reviews of these submissions are completed, they will be forwarded to FEMA for the final review and approval.

All of the work necessary to allow the State Fire Marshal to inspect the Dental School for re-occupancy is substantially complete. LSUHSC Facilities staff members have begun making their own tests of critical systems in preparation for the Fire Marshal’s visit, which will be scheduled for late next week. The final environmental remediation and duct cleaning at the Dental School is complete. All of the ceiling grid and tile has been installed on the upper floors of both the Clinic and Administration Buildings. The Central Sterilization project is proceeding on schedule. Work on the repairs to the exterior parking lot lighting and parking lot controls is on schedule and should be completed by the end of next week.

Bids were received for the ceiling work to be done on the 1st floor of the Clinic Building and an award is pending. Work is well underway to construct some temporary work space for the Facility Services staff, the Supply Chain Management staff, and the Dental School’s Central Supply function and progressing rapidly.

Temporary repairs continue to be made at 1542 Tulane Ave. The final stages of work to restore power to the remaining portions of the building are nearing completion. Power has been restored to the L'M parking lot guard house. The replacement of the flood damaged and obsolete fire alarm system is progressing well. The temporary emergency generator is on site and a change order has been proposed for the construction of a foundation on which it can be mounted. The air compressor system that is required by the building system controls is being replaced. The repairs to the building hot water heaters are nearing completion. Repairs are moving forward on a three additional elevators in the building.

Here are highlights of work that has been accomplished since our last update was posted two weeks ago.

Project to Make Temporary Repairs to the Dental School:

  • The final environmental remediation and duct cleaning contract is substantially complete.
    • Air handlers 6A, 8A, 8C and 3B in the Administration Building have passed the viable sampling tests and have been turned on.
  • The project to upgrade the network infrastructure is substantially complete.
    • The testing of all the new network cable installed in the Clinic Building has been completed.
    • The testing of all the new network cable installed in the Administration Building has been completed.
    • The construction of the new network/telecom closets is nearing completion.
      • There is still some electrical, lighting and HVAC work to be completed in the new network/telecom closets.
      • The contractor has not received the doors yet.
    • Firetrol is turning the sprinkler heads in the new network/telecom closets due to the fact that ceilings will not be installed in the closets.
  • The contract to replace the ceiling tiles and grid in the upper floors of the Clinic and Administration Building is complete.
    • The architects prepared a punch list for this project and provided it to the contractor.
    • The contractor has completed the punch list.
  • The contract to restore the Central Sterilization function in the 3rd floor location in the Clinic Building space is in place and the work is well underway.
    • The wall framing is complete.
    • The drywall installation is complete and the walls are taped and joint compound has been applied.
    • Fixtures, outlets and utilities have been roughed in.
    • The door frames and doors have been installed.
    • The painting of the walls in the Central Sterilization area is complete.
    • The ceiling grid has been installed.
    • New ductwork is connected and the installation of the diffusers has been completed.
    • The ceiling tiles have been installed.
  • The repairs to the exterior site lighting and the electrical work necessary to power the parking lot controls are continuing.
    • Approximately one half of the light poles and fixtures have been delivered and are being installed.
    • The remainder will be delivered and installed by the end of next week.
  • All of door hardware for the replacement doors and frames needed for the 1st floors has been delivered and the installation is complete.
  • Carpet and floor tile is being installed on the 2nd and 3rd floors of the Administration Building.
    • The installation of carpet in the Library is complete
    • The 2nd floor carpet and floor tile installation is nearing completion.
      • Tile has been laid in the 2nd floor Administration Building hallway near the clinic area.
    • The 4th floor Auditorium C is being prepped for the carpet installation and the re-carpeting of Auditorium C has begun.
  • A change order proposal to replace a number of HVAC inline dampers in the buildings has been received and approved.
    • LSUHSC had a limited number of dampers on site and several of the dampers have been installed.
    • Some of the old dampers have been found to be sealed with tape that contains asbestos. A small contract has been issued for the removal of the asbestos tape so that the remainder of the old dampers can be removed.
    • The rest of the dampers are on order and are expected any day.
  • A change order proposal to make repairs to the compressed air line running through the utility tunnel from the Power Plant to the Clinic Building and the related equipment has been received and approved.
    • Work is underway on these repairs and should be completed by early next week.
  • Representatives of the roof testing company completed their analysis and provided the results to the architects.
    • The test results and the architects’ recommendation to replace the roofs on the Clinic and Administration Building have been forwarded to FEMA for their consideration.
    • The architects are nearing completion of the construction and bid documents for this work. They are scheduled to be delivered early next week.
    • If FEMA concurs with the recommendation to replace the roofs, bids for the project can be advertised.
  • Schindler Elevator has been working on completing repairs to the three passenger elevators in the Administration Building.
  • Work on the Change Order to create temporary work spaces for Supply Chain Management, Facility Services and Central Supply on the 1st floor of the Clinic Building has begun.
    • Electricians are on site and started work.
      • The rough in for the new lights, electrical outlets, etc. is substantially complete.
    • Temporary walls have been extended to the ceiling.
    • Demolition of old duct work, electrical wiring, etc. is complete.
    • Temporary lighting has been installed.
    • The air handling units have been disconnected and moved to their temporary locations and connections have been completed.
    • The sanding of the terrazzo floor in the public areas of the 1st floor has begun in preparation for the re-finishing.
  • A number of water isolation valves have been installed throughout the Dental School buildings.
  • The installation of a domestic water booster pump and a vacuum breaker completed.
  • The curbing at the base of the stairs near the pay parking lot has been repaired.
    • Other walking surfaces are being repaired as needed.
  • The Dental School fire pump system was successfully tested and was "green tagged" in preparation for the inspection by the State Fire Marshal. 
    • The emergency generator was successfully tested.

Project to Make Temporary Repairs to 1542 Tulane Ave.:

  • Ordes Electric installed the electrical panel covers for sub-panels PLL-4, PLL-6, ' PLL-7 which power the rear section (oldest) of the building.
    • Ordes completed the installation of panel covers for electrical panels PLL-4, PLL-6, ' PLL-7.
    • Ordes began preparing substation PL-1 for the installation of the current transformer and meter.
    • Ordes installed the new fuses for the 2nd leg of the 4th floor riser.
    • Ordes Electrical is installing the disconnects for the condensate pump.
  • Ordes has running conduit to the guard shack for the L'M Bldg parking lot.
    • The electrical power has been connected to the guard shack.
    • The lights and air conditioning have been energized.
  • Ordes is making preparations for the installation of the temporary emergency generator.
  • Siemens is still removing the obsolete fire alarm devices and the related wiring on the upper floors of the building.
    • All of the obsolete fire alarm devices on Floors 1 through 8.
    • They have completed the installation of the new fire alarm cabinet in the control room.
    • New fire alarm panels are being installed in the 1st floor Security Office.
    • Cable is being pulled to the new panels.
  • Stratos continues their repair work on Elevator #2; installing the safeties, new rollers, guides, and the header for the cab doors.
    • They have installed the cab door operator and the cab doors.
    • The governor sheave and cables have been installed.
    • The travel cables have been installed.
    • The telephone line for Elevator #2 has been run to the elevator control room.
  • Stratos is also working on the installation of the controllers for Elevators #7 and #8, installing new conduit and wiring connections.
  • Arc Mechanical delivered some additional piping material to the site and continues to fabricate the piping for the hot water heaters in the basement.
    • Arc Mechanical installed the piping connections and assemblies (which include the valves and actuators) for the hot water heaters.
    • The hot water heater installation is nearing completion.

Downtown Campus Landscaping Project:

  • Crews completed the installation of irrigation valve wiring on the S. Roman St. and the Perdido St. sides of MEB.
  • The crews completed the installation the 2 inch irrigation main on the S. Roman St. side of the MEB and around the Perdido St. side to the corner with Bolivar St.
  • Crews installed irrigation valves on the S. Roman St. and Perdido St. sides of the MEB.
  • Valve boxes were installed on the S. Roman St. side of the Allied Health/Nursing Building.
  • The last of the trenches needed on the Perdido St. side and the Bolivar St. side of the MEB were finished.
  • Work crews completed digging trenches for the installation of irrigation wires and the 2 inch main water pipe to the Gravier St. side of the Allied Health/Nursing Building. 
    • They continued sealing around the new pipe and electrical wires at the ends of the sleeves installed under the sidewalks on the S. Roman St. and Perdido St. sides of the MEB.
  • Crews completed the installation of the short sleeves on the Bolivar St. and Gravier St. sides of the MEB and the Allied Health/Nursing Building.
  • Crews continued digging and preparing the trenches along the Bolivar St. side of the MEB and the Allied Health/Nursing Building where the irrigation control wire will be installed.

MEB Cafeteria Project:

  • The dry wall has been installed and the joints floated and sanded.
  • The installation of the ceiling material has been completed.
  • The ceiling light fixtures have been installed.
  • The sanding and finishing of the terrazzo flooring is finished.
  • The new air handling unit serving the dining area and the serving line was successfully started up.

CSRB 1st Floor Vivarium:

  • Arc Mechanical completed the process of extending the piping in preparation for elevating the Animal Care chiller.

MEB Duct Cleaning:

  • The 4th floor cleaning is complete.
  • The 3rd floor cleaning is complete.
  • The 2nd floor cleaning is complete.
  • The cleaning has begun on the 5th floor.

Campus Security:

  • One of the initiatives to enhance the security of the campus is to improve on the exterior lighting. A project is underway to replace all of the old lights in the Roman St. Garage with new lights that are both brighter and more energy efficient.
    • The result will be brighter light throughout the parking garage, without any increase in energy consumption.
    • Lights have been replaced on the majority of the 4th, 5th and 6th floors of the garage.
      • The crew began working at the top of the structure and they are working their way down.
  • Parking Lot #2:
    • The contractors have installed three new light poles and fixtures in the parking lot.

We encourage you to follow the details of our progress through the daily updates which are posted on the web site.